Write a balanced chemical equation for the reaction that occurs when:
-Magnesium metal reacts with chlorine gas.
-Barium carbonate decomposes into barium oxide and carbon dioxide when heated.
-CH3OCH3 gas is combusted in air.
1
Expert's answer
2016-12-11T05:09:08-0500
- Mg (s) + 2HCl (g) = MgCl2 (s) + H2 (g)
- BaCO3 (s) = BaO (s) + CO2 (g)
- CH3OCH3 (g) + 3O2 (g) = 2CO2 (g) + 3H2O (g)